OpenStudy (anonymous):

Suppose that r varies directly as p and inversely as q2, and that r = 27 when p = 3 and q = 2. Find r when p = 2 and q = 3.

OpenStudy (anonymous):

q^2, r=kp/q^2 Need to find "k" Since r=27 when p=3 and q=2, 27=k(3)/2^2 27=3k/4 108=3k k=36 Equation: r=(36)p/q^2 Now find r when p=2 and q=3. r=(36)2/9 r=8